OpenStudy (anonymous):

the following sequence is an arithmetic sequence.you are now shown very many terms of this sequence,but you are shown enough to fill in the missing number.what number goes in the blank? 14, _,56

OpenStudy (chaise):

The answer is 28. Do you need further help on how this solution was made?

OpenStudy (anonymous):

no thanks

OpenStudy (dumbcow):

i disagree, because it is arithmetic not geometric 14,x,56 An = An-1 + d x = 14 +d --> d = x-14 56 = x +d --> d = 56-x By substitution x-14 = 56-x 2x = 70 x = 35 the sequence is 14,35,56